# Long-Tailed 3D Detection via Multi-Modal Late-Fusion

### v4

@@ 3 LT3D Benchmarking Protocol
Conceptually, LT3D extends the traditional 3D detection problem, which focuses on identifying objects from $K$ common classes, by further requiring detection of $N$ rare classes. As LT3D emphasizes detection performance on all classes, we report metrics for three groups of classes based on their cardinality (Fig. 1-left): many ($>$50k instance/class), medium (5$\sim$50k instance/class), and few ($<$5k instance/class). To better analyze LT3D performance, we present two metrics below.
@@
Hierarchical Mean Average Precision (mAPH). For safety-critical applications, it is more desirable to localize-but-misclassify an object than miss-detect this object, e.g., detecting but misclassifying a child as adult is better than not detecting this child. Therefore, we present hierarchical mean average precision (mAPH), which considers such semantic relationships across classes to award partial credit. We use APH as a supplementary diagnostic tool to analyze how LT3D detectors make mistakes. To encode inter-class relationships, we leverage the semantic hierarchy (Fig. 4) such as the one defined by nuScenes [6]. We derive partial credit as a function of semantic similarity using the least common ancestor (LCA) distance metric. LCA has been proposed for image classification [56, 57] but not for object detection. Extending this metric to object detection is challenging as we must jointly evaluate semantic and spatial overlap. For clarity, we describe how to compute APH for a class $C$.
@@
4 LT3D Methods: The Devil Is In The Details
To address LT3D, we first retrain state-of-the-art (SOTA) detectors on all long-tailed classes. We consider the following detectors:
@@ 4 LT3D Methods: The Devil Is In The Details
  • End-to-End Multi-modal 3D Detectors: TransFusion [34], BEVFusion [35], CMT [10], and DeepInteraction [59].
In Section 4.1 and 4.2, we introduce two modifications that consistently improve LT3D performance, and describe our multi-modal late-fusion framework.
Figure 4: The nuScenes dataset defines a semantic hierarchy for all annotated classes (Fig. 1). We highlight common classes in white and rare classes in gold. The standard nuScenes benchmark makes two choices for dealing with rare classes: (1) ignore them (e.g., stroller and pushable-pullable), or (2) group them into coarse-grained classes (e.g., adult, child, construction-worker, police-officer are grouped as pedestrian). Since the pedestrian class is dominated by adult (Fig. 1), the standard benchmarking protocol masks the challenge of detecting rare classes like child and police-officer. We explicitly incorporate this semantic hierarchy to evaluate LT3D performance via the proposed mAPH metric described in Sec. 3.
@@ 4.1 Grouping-Free Detector Head
Extending existing 3D detectors to train with more classes is surprisingly challenging. Many contemporary networks use a multi-head architecture that groups classes of similar size and shape to facilitate efficient feature sharing [18, 12]. For example, CenterPoint [12] groups pedestrian and traffic-cone since these objects are both tall and skinny. However, this multi-head grouping design may not work for super-classes like pushable-pullable and debris that contain diverse objects of different sizes and shapes. Moreover, in the multi-head design, each head is a group-specific detector that consists of several layers with lots of parameters. This makes multi-head grouping difficult to scale for a large number of classes. To address these issues, we treat each class as its own group to avoid hand-crafted grouping heuristics and design a group-free strategy, in which each class has only one linear layer as its detector (Fig. 7) and all classes share a single detector head. This design significantly reduces the number of parameters and allows learning the shared feature backbone collaboratively with all classes, effectively mitigating overfitting to rare classes. Adding a new class is as simple as adding a single linear layer to the detector head. Importantly, we demonstrate that our grouping-free detector head outperforms conventional grouping-based methods in the supplement.
@@ 4.2 Training with a Semantic Hierarchy
The nuScenes dataset defines a semantic hierarchy (Fig. 4) for all classes, grouping semantically similar classes under coarse-grained categories [6]. We leverage this hierarchy during training. Specifically, we train detectors to predict three labels for each object: its fine-grained class (e.g., child), its coarse class (e.g., pedestrian), and the root class object. We adopt a grouping-free detector head (Fig. 7) that outputs separate “multitask” heatmaps for each class, and use a per-class sigmoid focal loss [60] rather than multi-class cross-entropy loss to avoid normalizing class probabilities across fine-grained and super-classes. It is worth noting that our approach does not explicitly enforce a tree-like hierarchy, and can be applied to more complex label relationships [61]. Crucially, adding a vehicle heatmap does not directly interfere with the car heatmap. However, this might produce repeated detections on the same test object (e.g. a single ground-truth car may be detected as a car, vehicle, and object). We address that by simply ignoring coarse detections at test time. We explore alternatives in the supplement and conclude that they achieve similar LT3D performance. Perhaps surprisingly, this training method improves detection performance not only for rare classes, but also for common classes (Table IV).
4.3 Multimodal Filtering for Detection Fusion
Rare class instances are often small and can be challenging to recognize from sparse (LiDAR) geometry alone – even humans struggle to find strollers in LiDAR point clouds. This suggests that one can leverage multi-modal cues to improve LT3D. First, we find that, although LiDAR-based detectors are widely adopted for 3D detection, they produce many high-scoring false positives (FPs) for rare classes due to misclassification. We focus on removing such FPs. To this end, we use an RGB-based 3D detector (e.g., FCOS3D [13]) to filter out high-scoring false-positive LiDAR detections by leveraging two insights: (1) LiDAR-based 3D-detectors can achieve high recall and precise 3D localizations for TPs, and (2) RGB-based 3D-detections are accurate w.r.t recognition although their 3D localization is poor. Fig. 5 demonstrates our Multi-Modal Filtering (MMF) strategy. For each RGB detection, we keep LiDAR detections within a radius of $m$ meters and remove all the other LiDAR detections. We denote this method as MMF($D_{L}$, $D_{R}$), where DL and $D_{R}$ are any 3D LiDAR detector and 3D RGB detector, respectively. Table II demonstrates that MMF greatly improves LT3D.
@@ 4.3 Multimodal Filtering for Detection Fusion
Figure 5: Multi-Modal Filtering (MMF) effectively removes high-scoring false-positive LiDAR detections. The green boxes are ground-truth strollers, while the blue boxes are stroller detections from the LiDAR-based detector CenterPoint [12] (left) and RGB-based detector FCOS3D [13] (mid). The final filtered result removes LiDAR detections not within $m$ meters of any RGB detection, shown in the red region, and keeps all other LiDAR detections, shown in the white region (right).
@@
Refer to caption Refer to caption
@@
Figure 6: We analyze our multi-modal filtering (MMF) approach. Rare classes are most often confused by the dominant class (in blue) in each superclass: (left) Vehicle is dominated by car, (mid) Pedestrian is dominated by adult, and (right) Movable is dominated by barrier. We find that class confusions are reasonable. Car is often mistaken for truck. Similarly, truck, construction-vehicle and emergency-vehicle are most often mistaken for car. Bicycle and motorcycle are sometimes misclassified as car, presumably because they are sometimes spatially close to cars. Adults have similar appearance to police-officer and construction-worker, and they are often co-localized with child and stroller.
Limitations of Multi-Modal Filtering. Despite the effectiveness of multi-modal filtering, this approach is sensitive to the classification accuracy of FCOS3D when matching LiDAR and RGB-based detections, leading to many correctly localized but misclassified detections. We use a confusion matrix to further analyze the misclassifications within each superclass, as shown in Fig. 6. We explain how to compute a confusion matrix for the detection task: For each superclass, we construct a confusion matrix, in which the entry $(i,j)$ indicates the misclassification rate of class-$i$ objects as class-$j$. Specifically, given a class $i$, we find its predictions that match ground-truth boxes within a 2m center-distance of class-$i$ and all its sibling classes (LCA=1, within the corresponding superclass); we ignore all unmatched detections. This allows us to count the mis-classifications of class-$i$ objects into class-$j$. We normalize these counts to produces misclassification rates. We find that rare classes are most often confused by the dominant class in each superclass: Vehicle is dominated by car, Pedestrian is dominated by adult, and Movable is dominated by barrier. We posit that this can be addressed with more accurate RGB-based classifiers, motivating our study of multi-modal late-fusion (MMLF) from first principles.
4.4 Delving into Multi-Modal Late-Fusion for LT3D
@@
Motivated by the effectiveness of our frustratingly simple multi-modal filtering (MMF) approach, we delve into this multi-modal late-fusion (MMLF) strategy. We explore three questions from first principles (Fig. 3): (1) how do we effectively incorporate RGB information, (2) how do we match RGB and LiDAR detections, and (3) how do we fuse them. We denote our final approach as MMLF($D_{L}$, $D_{R}$) in Table I, where DL is any 3D LiDAR detector and $D_{R}$ is any 2D RGB detector.
4.4.1 How Do We Incorporate RGB Information?
@@
Although LiDAR offers accurate localization, LiDAR-only detectors struggle to recognize objects using sparse LiDAR alone. RGB images provide complementary information that is essential for identifying objects and disambiguating those that are geometrically similar in point clouds but semantically different in images, e.g., construction worker vs. police officer. In the last subsection, we proposed multimodal filtering (MMF) (Fig. 5), a strategy that ensembles a 3D RGB detector and a 3D LiDAR detector, yielding remarkable improvements for LT3D (Table II). However, we find that ensembling a 2D RGB detector and a 3D LiDAR detector yields more significantly better LT3D performance. We present insights on why using a 2D detector to incorporate RGB information is better and ablate the impact of using 2D vs. 3D RGB detectors for late-fusion in Table III.
@@
2D RGB detectors are more mature. 2D object detection is a fundamental problem in computer vision [62, 55, 63] that has matured in recent years and model trade-offs are well understood [63, 15, 64, 60]. In this work, we consider two state-of-the-art 2D RGB detectors, YOLOV7 [58] and DINO [11]. YOLOV7 is a real-time detector that identifies a number of training techniques that nearly doubles the inference efficiency over prior work without sacrificing performance. DINO [11] is a recent transformer-based detector that improves upon DETR [16] using denoising anchor boxes. As 2D detectors do not make 3D predictions (e.g., depth and rotation), understanding how to best leverage them in the context of LT3D is a key challenge.
@@
Figure 8: Our multi-modal late-fusion (MMLF) strategy takes 3D LiDAR and 2D RGB detections as input (Fig. 3A), matches 2D RGB and (projected) 3D LiDAR detections on the image plane (Fig. 3B), and fuses matched predictions with score calibration and probabilistic ensembling to produce 3D detections (Fig. 3C).
4.4.2 How Do We Match Uni-Modal Detections?
Finding correspondence between two sets of RGB and LiDAR uni-modal detections is an essential step in the late-fusion framework (Fig. 3B). Our previous multi-modal filtering method [9] matched 3D RGB and 3D LiDAR detections using center distance (Fig. 5). However, precisely matching detections in 3D is difficult due to depth estimation errors from 3D RGB detectors. Instead, we opt to match 2D RGB and 3D LiDAR detections. Prior work attempts to inflate 2D detections to 3D (using LiDAR points [70]), but we find that matching inflated 2D detections also yields noisy 3D boxes (due to outlier points from the background) and reduces overall match quality. In contrast, we match uni-modal detections by projecting 3D LiDAR detections onto the 2D image plane, avoiding additional noise due to imprecise depth estimates. We ablate the impact of matching in 3D versus on the 2D image plane in Table III, and present our 2D matching algorithm below.
Spatial matching on the 2D image plane. Using the available sensor extrinstics, we project 3D LiDAR detections onto the 2D image plane. We then use the IoU metric to determine overlap between (projected) LiDAR and 2D RGB detections. A 2D RGB detection and a (projected) 3D LiDAR detection are considered a match if their IoU is greater than a threshold. Although conceptually simple, this matching method works significantly better than using center distance to match detections in 3D (cf. bottom two rows in Table III). Spatially matching uni-modal detections using 2D IoU yields three categories of detections: matched detections, unmatched RGB detections (that do not have corresponding LiDAR detections), and unmatched LiDAR detections (that do not have corresponding RGB detections). We handle unmatched detections below and present our fusion strategy for matched detections in the next subsection.
Handling unmatched detections. We remove unmatched 2D RGB detections, positing that any unmatched RGB detections are likely to be false positives given that LiDAR detectors tend to yield high recall [9]. Further, accurately inflating 2D RGB detections to 3D remains challenging. For unmatched 3D LiDAR detections, we down-weight their detection confidence scores by $w$ (inspired by SoftNMS [71]). We optimize for $w$ by grid search on a validation set and set $w=0.4$.
4.4.3 How Do We Fuse Matched Uni-Modal Detections?
@@
Addressing semantic disagreement between modalities. If the two modalities predict different semantic classes, we use the confidence score (which is calibrated as explained below) and class label of the RGB-based detection and the 3D box extent from the LiDAR-based detection. Intuitively, RGB detectors can predict semantics more reliably from high resolution images than LiDAR-only detectors. This helps correct misclassifications of geometrically similar but semantically different objects produced by the 3D LiDAR detector, as shown in Fig. 9. Importantly, prior late-fusion methods like CLOCS [38] and our MMF only perform late-fusion on matched predictions with semantic agreement and do not fix misclassifications. We find that handling such misclassifications is critical for improving rare class performance (Table I).
If both modalities predict the same semantic class, we perform score fusion and probablistic ensembling [72] as described below. Note that the confidence scores of RGB and LiDAR detections are not directly comparable by default: LiDAR-based detectors are often under-confident as it is difficult to distinguish foreground-vs-background using sparse LiDAR alone. Therefore, score calibration is crucial for fusion. Below, we explore score calibration of RGB ($x_{\text{RGB}}$) and LiDAR ($x_{\text{LiDAR}}$) detections.
Score calibration. We calibrate detection confidences per model by tuning a temperature $\tau_{c}$ for the logit score of class $c$ on a validation set before applying a sigmoid transform [73, 72], i.e., sigmoid(logitc/$\tau_{c}$). Optimally tuning per-class $\tau_{c}$ is computationally expensive as it requires tuning for all classes at the same time. Instead, we choose to greedily tune each $\tau_{c}$, optimizing per-class AP on a val-set for each class progressively ordered by their cardinalities. It is worth noting that this score calibration is only performed once in training and tuned $\tau_{c}$ and $p(c)$ do not need further optimization during inference. Importantly, score calibration does not increase runtime or complexity.
Probabilistic ensembling. We assume independent class prior $p(c)$ and conditional independence given the class label $c$ [72], i.e., $p(x_{\text{RGB}},x_{\text{LiDAR}}|c)=p(x_{\text{RGB}}|c)p(x_{\text{LiDAR}}|c)$. We compute the final score as
$\small p(c|x_{\text{RGB}},x_{\text{LiDAR}})$ $=p(x_{\text{RGB}},x_{\text{LiDAR}}|c)p(c)\ \ /\ \ p(x_{\text{RGB}},x_{\text{LiDAR}})$
@@
$\propto\ \ p(c|x_{\text{RGB}})p(c|x_{\text{LiDAR}})\ \ /\ \ {p(c)}$
where $p(c|x_{\text{RGB}})$ and $p(c|x_{\text{LiDAR}})$ are the posteriors after calibration. Unlike the balanced class distribution studied in [72], $p(c)$ can significantly impact the final long-tailed performance. To maximize detection accuracy, the optimal tuning practice is to jointly optimize all class priors $p(c)$ altogether, which is computationally expensive. Similar to score calibration, we greedily tune $p(c)$, one by one ordered by class cardinality. Probabilistic ensembling does not increase inference time or complexity.

### v5

@@ 3 LT3D Benchmarking Protocol
[…]
@@
Hierarchical Mean Average Precision (mAPH). For safety-critical applications, it is more desirable to localize-but-misclassify an object than fail to detect this object, e.g., detecting but misclassifying a child as adult is better than not detecting this child. Therefore, we propose hierarchical mean average precision (mAPH), which considers such semantic relationships across classes to award partial credit. We use APH as a supplementary diagnostic tool to analyze how LT3D detectors make mistakes. To encode inter-class relationships, we leverage the semantic hierarchy (Fig. 4) such as the one defined by nuScenes [5]. We derive partial credit as a function of semantic similarity using the least common ancestor (LCA) distance metric. LCA has been proposed for image classification [53, 62] but not for object detection. Extending this metric to detection is non-trivial as we must jointly evaluate semantic and spatial overlap. Below, we explain how to compute APH for a class $C$.
@@
Figure 4: The nuScenes dataset defines a hierarchy for the annotated classes (Fig. 1). We mark common classes in white and rare classes in gold. The standard nuScenes benchmark makes two choices for dealing with rare classes: (1) ignoring them (e.g., stroller and pushable-pullable), or (2) grouping them into super classes (e.g., adult, child, construction-worker, police-officer are grouped as pedestrian). As the pedestrian class is dominated by adult (Fig. 1), the standard benchmarking protocol masks the challenge of detecting rare classes like child and police-officer. We explicitly incorporate this hierarchy to evaluate LT3D performance via the proposed mAPH metric described in Sec. 3.
4 Our Proposed Approach
To address LT3D, we first retrain well-established detectors on all long-tailed classes. We consider the following detectors:
@@ 4 Our Proposed Approach
  • End-to-End Multi-modal 3D Detectors: TransFusion [2], BEVFusion [40], CMT [80], DeepInteraction [79], IS-Fusion [81], and EA-LSS [28].
In Section 4.1 and 4.2, we introduce two modifications that consistently improve LT3D performance. In Section 4.3, we present a simple yet effective late-fusion method.
@@ 4.1 Grouping-Free Detector Head
Extending existing 3D detectors to train with more classes is surprisingly challenging. Many contemporary networks use a multi-head architecture that groups classes of similar size and shape to facilitate efficient feature sharing [84, 82]. For example, CenterPoint [82] groups pedestrian and traffic-cone since these objects are both tall and skinny. This multi-head grouping design may not work for super-classes like pushable-pullable and debris that contain diverse objects of different sizes and shapes. Moreover, in the multi-head design, each head is a group-specific detector that consists of several layers with lots of parameters. This makes multi-head grouping difficult to scale for a large number of classes. To address these issues, we treat each class as its own group to avoid hand-crafted grouping heuristics and design a group-free strategy, in which each class has only one linear layer as its detector (Fig. 5) and all classes share a single detector head. This design significantly reduces the number of parameters and allows learning the shared feature backbone collaboratively with all classes, effectively mitigating overfitting to rare classes. Adding a new class is as simple as adding a single linear layer to the detector head. Importantly, we show that our grouping-free detector outperforms conventional grouping-based methods in Appendix F.
@@ 4.2 Training with a Semantic Hierarchy
The nuScenes dataset defines a semantic hierarchy (Fig. 4) that groups semantically similar classes under coarse-grained categories [5]. We leverage this hierarchy during training. Specifically, we train detectors to predict three labels for each object: its fine-grained class (e.g., child), its coarse class (e.g., pedestrian), and the root class object. We adopt a grouping-free detector head (Fig. 5) that outputs separate “multitask” heatmaps for each class, and use a per-class sigmoid focal loss [35] rather than multi-class cross-entropy loss to avoid normalizing class probabilities across fine-grained and super-classes. It is worth noting that our approach does not explicitly enforce a tree-like hierarchy, and can be applied to more complex label relationships [39]. Crucially, adding a vehicle heatmap does not directly interfere with the car heatmap. However, this might produce repeated detections on the same test object (e.g. a single ground-truth car may be detected as a car, vehicle, and object). We address that by simply ignoring coarse detections at test time. We explore alternatives in Appendix G and conclude that they achieve similar LT3D performance. Perhaps surprisingly, this training method improves detection performance not only for rare classes, but also for common classes (Table 4).
4.3 Multi-Modal Filtering for Detection Fusion
Rare-class instances are often small and can be challenging to recognize from sparse (LiDAR) geometry alone: even humans struggle to find strollers in LiDAR point clouds. This suggests that one can leverage multi-modal cues to improve LT3D. First, we find that, although LiDAR-based detectors are widely adopted for 3D detection, they produce many high-scoring false positives (FPs) for rare classes due to misclassification. We focus on removing such FPs. To this end, we use an RGB-based 3D detector (e.g., FCOS3D [76]) to filter out high-scoring false-positive LiDAR detections by leveraging two insights: (1) LiDAR-based 3D-detectors can achieve high recall and precise 3D localizations for TPs, and (2) RGB-based 3D-detections are accurate w.r.t recognition although their 3D localization is poor. Fig. 6 demonstrates our Multi-Modal Filtering (MMF) strategy. For each RGB detection, we keep LiDAR detections within a radius of $m$ meters and remove all the other LiDAR detections. We denote this method as MMF($D_{L}$, $D_{R}$), where DL and $D_{R}$ are any 3D LiDAR detector and 3D RGB detector, respectively. Table 2 demonstrates that MMF greatly improves LT3D.
@@ 4.3 Multi-Modal Filtering for Detection Fusion
Figure 6: The proposed method Multi-Modal Filtering (MMF) effectively removes high-scoring false-positive LiDAR detections. The green boxes are ground-truth strollers, while the blue boxes are stroller detections from the LiDAR-based detector CenterPoint [82] (left) and RGB-based detector FCOS3D [76] (mid). The final filtered result removes LiDAR detections not within $m$ meters of any RGB detection, shown in the red region, and keeps all other LiDAR detections, shown in the white region (right).
Limitations of MMF. Despite the effectiveness of multi-modal filtering, our approach is sensitive to the classification accuracy of FCOS3D when matching LiDAR and RGB-based detections, leading to many correctly localized but misclassified detections. We use a confusion matrix to further analyze the misclassifications within each superclass, as shown in Fig. 7. We explain how to compute a confusion matrix for the detection task: For each superclass, we construct a confusion matrix, in which the entry $(i,j)$ indicates the misclassification rate of class-$i$ objects as class-$j$. Specifically, given a class $i$, we find its predictions that match ground-truth boxes within a 2m center-distance of class-$i$ and all its sibling classes (LCA=1, within the corresponding superclass); we ignore all unmatched detections. This allows us to count the mis-classifications of class-$i$ objects into class-$j$. We normalize these counts to produces misclassification rates. We find that rare classes are most often confused by the dominant class in each superclass: Vehicle is dominated by car, Pedestrian is dominated by adult, and Movable is dominated by barrier. We posit that this can be addressed with more accurate RGB-based classifiers, motivating our study of multi-modal late-fusion (MMLF) from first principles, presented next.
@@
Refer to caption Refer to caption Refer to caption
@@
Figure 7: We analyze our multi-modal filtering (MMF) approach. Rare classes are most often confused by the common class (in blue) in each superclass: (left) Vehicle is dominated by car, (mid) Pedestrian is dominated by adult, and (right) Movable is dominated by barrier. We find that class confusions are reasonable. Car is often mistaken for truck. Similarly, truck, construction-vehicle and emergency-vehicle are most often mistaken for car. Bicycle and motorcycle are sometimes misclassified as car, presumably because they are sometimes spatially close to cars. Adults have similar appearance to police-officer and construction-worker, and they are often co-localized with child and stroller.
5 Multi-Modal Late-Fusion for LT3D
@@
Motivated by the effectiveness of our frustratingly simple multi-modal filtering (MMF) approach, we investigate this multi-modal late-fusion (MMLF) strategy further. We explore three questions from first principles (Fig. 3): (1) how to effectively incorporate RGB information, (2) how to match RGB and LiDAR detections, and (3) how to fuse them.
5.1 How Do We Incorporate RGB Information?
@@
Although LiDAR offers accurate localization, LiDAR-only detectors struggle to recognize objects using sparse LiDAR alone. RGB images provide complementary information that is essential for identifying objects and disambiguating those that are geometrically similar in point clouds but semantically different in images, e.g., construction worker vs. police officer. In the last subsection, we proposed multimodal filtering (MMF) (Fig. 6), a simple late-fusion strategy that ensembles a 3D RGB detector and a 3D LiDAR detector and yields remarkable improvements for LT3D (Table 2). Motivated by the late-fusion strategy, we further explore fusing a 2D RGB detector and a 3D LiDAR detector. Surprisingly, this yields more significantly better LT3D performance. We present insights on why using a 2D detector to incorporate RGB information is better and ablate the impact of using 2D vs. 3D RGB detectors for late-fusion in Table 3.
Figure 8: Our multi-modal late-fusion (MMLF) strategy takes 3D LiDAR and 2D RGB detections as input (Fig. 3A), matches the detections on the image plane (Fig. 3B), and fuses matched ones with score calibration and probabilistic fusion to produce the final 3D detections (Fig. 3C).
@@
2D RGB detectors are more mature. 2D object detection is a fundamental problem in computer vision [16, 36, 57] that has matured in recent years and model trade-offs are well understood [57, 33, 52, 35]. Our work considers two well-established 2D RGB detectors, YOLOV7 [70] and DINO [88]. YOLOV7 is a real-time detector that identifies a number of training techniques that nearly doubles the inference efficiency over prior work without sacrificing performance. DINO is a transformer-based detector that improves upon DETR [8] using denoising anchor boxes. As 2D detectors do not estimate 3D attributes (e.g., depth and rotation), understanding how to best leverage them in the context of LT3D is a key challenge (explored in Section 5.2).
@@
5.2 How Do We Match Uni-Modal Detections?
Finding correspondence between two sets of RGB and LiDAR uni-modal detections is an essential step in the late-fusion framework (Fig. 3B). Our previous MMF method matches 3D RGB and 3D LiDAR detections using center distance in the 3D space (Fig. 6). However, precisely matching them in the 3D space is error-prone due to depth estimation errors from the 3D RGB detector. Hence, we opt to match 2D RGB (without the need to estimate depth in detection) and 3D LiDAR detections. Prior work attempts to inflate 2D detections to 3D for matching by using LiDAR points [72], but we find that matching inflated 2D detections also yields noisy 3D boxes (due to outlier LiDAR points from the background) and reduces overall match quality. In contrast, we match uni-modal detections by projecting 3D LiDAR detections onto the 2D image plane, avoiding additional noise due to imprecise depth estimates. We ablate the impact of matching in 3D versus on the 2D image plane in Table 3, and present our 2D matching algorithm below.
Spatial matching on the 2D image plane. Using the available sensor extrinstics, we project 3D LiDAR detections onto the 2D image plane. We then use the IoU metric to determine overlap between projected LiDAR and 2D RGB detections. A 2D RGB detection and a projected 3D LiDAR detection are considered a match if their IoU is greater than a threshold. Although conceptually simple, this matching method works significantly better than using center distance to match detections in 3D (cf. bottom two rows in Table 3). Spatially matching uni-modal detections using 2D IoU yields three categories of detections: matched detections, unmatched RGB detections (that do not have corresponding LiDAR detections), and unmatched LiDAR detections (that do not have corresponding RGB detections). We handle unmatched detections below and show how to fuse matched ones in the next subsection.
Handling unmatched detections. We remove unmatched 2D RGB detections, positing that any unmatched RGB detections are likely to be false positives given that LiDAR detectors tend to yield high recall [46] (Appendix M). For unmatched 3D LiDAR detections, we down-weight their detection confidence scores by $w$ (inspired by SoftNMS [3]). We optimize for $w$ by grid search on a validation set and set $w=0.4$.
5.3 How Do We Fuse Matched Uni-Modal Detections?
@@
Addressing semantic disagreement between modalities. If the two modalities predict different semantic classes, we use the confidence score (which is calibrated as explained below) and class label of the RGB-based detection, and the 3D box from the LiDAR-based detection. Intuitively, RGB detectors can predict semantics more reliably from high resolution images than LiDAR-only detectors. This helps correct misclassifications of geometrically similar but semantically different objects produced by the 3D LiDAR detector, as shown in Fig. 9. To note, prior late-fusion methods like CLOCS [49] and our MMF [46] only fuse matched predictions with semantic agreement and do not correct misclassifications. In contrast, ours handles misclassifications that significantly improves rare class performance (Table 1).
Fusing semantically-agreed detections. If both modalities predict the same semantic class, we perform score fusion and probablistic fusion [9] as described below. Note that the confidence scores of RGB and LiDAR detections are not directly comparable: LiDAR-based detectors are often under-confident as it is difficult to distinguish foreground-vs-background using sparse LiDAR alone. Therefore, score calibration is crucial for fusion. Below, we explore score calibration of RGB ($x_{\text{RGB}}$) and LiDAR ($x_{\text{LiDAR}}$) detections.
[…]
Probabilistic fusion. We assume independent class prior $p(c)$ and conditional independence given the class label $c$ [9], i.e., $p(x_{\text{RGB}},x_{\text{LiDAR}}|c)=p(x_{\text{RGB}}|c)p(x_{\text{LiDAR}}|c)$. We compute the final score as
$\small p(c|x_{\text{RGB}},x_{\text{LiDAR}})$ $=\frac{p(x_{\text{RGB}},x_{\text{LiDAR}}|c)p(c)}{p(x_{\text{RGB}},x_{\text{LiDAR}})}$
@@
$\propto\ \ \frac{p(c|x_{\text{RGB}})p(c|x_{\text{LiDAR}})}{{p(c)}}$
where $p(c|x_{\text{RGB}})$ and $p(c|x_{\text{LiDAR}})$ are the posteriors after calibration. Unlike assuming balanced class prior distributions in [9], we tune $p(c)$ in a greedy fashion for classes sorted by their cardinality. Similar to temperature tuning in score calibration, probabilistic fusion does not increase inference time (Appendix E).
Method Modality All Many Medium Few
FCOS3D [76] (ICCVW’21) C 20.9 39.0 23.3 2.9
BEVFormer [43] (ECCV’22) C 27.3 52.3 31.6 1.4
PolarFormer [31] (AAAI’23) C 28.0 54.0 31.6 2.2
CenterPoint [82] (CVPR’21) L 40.4 77.1 45.1 4.3
TransFusion-L [2] (CVPR’22) L 38.5 68.5 42.8 8.4
BEVFusion-L [40] (ICRA’23) L 42.5 72.5 48.0 10.6
CMT-L [80] (ICCV’23) L 34.7 73.4 35.9 1.1
CLOCS [49] (IROS’20) C+L 40.0 68.2 45.7 10.0
TransFusion [2] (CVPR’22) C+L 39.8 73.9 41.2 9.8
DeepInteraction [79] (NeurIPS ’22) C+L 43.7 76.2 51.1 7.9
BEVFusion [40] (ICRA’23) C+L 45.5 75.5 52.0 12.8
CMT [80] (ICCV’23) C+L 44.4 79.9 53.0 4.8
EA-LSS [28] (arXiv’23) C+L 44.6 79.5 52.6 6.2
IS-Fusion [81] (CVPR’24) C+L 44.1 77.9 59.4 3.8
CenterPoint [82] + RCNN [18] C+L 34.0 64.8 37.5 4.3
MMF(CenterPoint, FCOS3D) C+L 43.6 77.1 49.0 9.4
MMLF(CenterPoint, DINO) C+L 51.4 77.9 59.4 20.0
Table 1: Benchmarking results on nuScenes. We denote the RGB and LiDAR modalities by C and L, respectively. Our simple late-fusion method MMF, which uses detections of the 3D RGB-based detector FCOS3D to filter detections of the 3D LiDAR-based detector CenterPoint, rivals prior methods. Furthermore, our final method MMLF, which fuses 3D LiDAR and 2D RGB detections (from CenterPoint and DINO, respectively) with score calibration and probabilistic fusion, performs the best averaged across all common and rare classes. Notably, it outperforms end-to-end multi-modal methods such as BEVFusion [40], DeepInteraction [79], and CMT [80]. Importantly, our MMLF nearly doubles the detection performance achieved by prior work on classes with few examples!